OpenStudy (anonymous):

If the area of a circle is 36π square feet, what is the radius, in feet?

OpenStudy (goldphenoix):

Do you know the formula to finding the area of a circle?

OpenStudy (anonymous):

a=πr^2

OpenStudy (goldphenoix):

Great. We know that the area of a circle is 36π. So we can plug that in our equation! \[\large \large r^2\pi = 36 \pi\] Make sense so far?

OpenStudy (anonymous):

yes

OpenStudy (goldphenoix):

Great! Now what's pi divided by pi give us?

OpenStudy (anonymous):

pi squared

OpenStudy (goldphenoix):

Not quite. pi multiply by pi give us pi squared. We're doing pi divided by pi. What do you think it is now? TIP: The 2 pi have the same value.

OpenStudy (anonymous):

its just pi

OpenStudy (goldphenoix):

Hmm. We know that pi and the other pi has the same value. So if we have the same value, divided by the same value, then what does it give us? Example: 10 / 10 = ? 2.10 / 2.10 = ?

OpenStudy (anonymous):

o so its 9π

OpenStudy (goldphenoix):

o.o? Nope. Solve my example please: 10 / 10 give us? 2.10 / 2.10 give us?

OpenStudy (anonymous):

they both 1

OpenStudy (goldphenoix):

Yes! Now any value divided by the same value will always equal 1, except special circumstances. So what do you think pi divided by pi give us?

OpenStudy (anonymous):

pi

OpenStudy (goldphenoix):

Nooo! Pi is about 3.14. So what's 3.14 divided by 3.14 give us?

OpenStudy (anonymous):

pi which = 3.14

OpenStudy (goldphenoix):

Pi is appx. 3.14. We just rounded, but it's still the same. So what's 3.14 / 3.14 give us?

OpenStudy (goldphenoix):

Fine. Lets round it even more. What's 3 / 3 give us?

OpenStudy (anonymous):

3.14/3.14=1

OpenStudy (goldphenoix):

Good! So pi divided by pi cancel out. Now we have: \[\large \large r^2 = 36\] Now how can we get r by itself?

OpenStudy (anonymous):

divide

OpenStudy (goldphenoix):

Divide what?

OpenStudy (anonymous):

by r

OpenStudy (goldphenoix):

Hmm. What will it give us?

OpenStudy (anonymous):

idk

OpenStudy (goldphenoix):

We can't divide. Do you know what's the opposite of square is?

OpenStudy (anonymous):

is the answer 6π

OpenStudy (goldphenoix):

What?!?!?! Nooo! Pi is gone!

OpenStudy (goldphenoix):

We have r^2 = 36 There's no pi in this equation, so we don't need to worry about pi. Have you ever heard of square root?

OpenStudy (anonymous):

so its either 6 and yes

OpenStudy (goldphenoix):

What's the square root of r^2?

OpenStudy (anonymous):

\[pir ^{2}=36\pi then r=6\]

OpenStudy (goldphenoix):

The answer is not just 6. But it's close. :)

OpenStudy (goldphenoix):

e.e.